OpenStudy (anonymous):

Antonio is buying new rims and tires for his SUV. The total cost of the rims and tires is $2300. If the cost of the tires is $200 more than twice as much as the rims, what is the cost of the rims?

OpenStudy (anonymous):

Use these two equations: t + r = 2300 (total cost is $2300 for both) t = 2r + 200 (the tires are $200 more than twice the cost of the rims) Substitute 2r + 200 (from equation 2) for t in equation 1: (2r + 200) + r = 2300 Combine like terms on the left side: 3r + 200 = 2300 Subtract 200 from both sides: 3r = 2100 Divide both sides by 3: r = 700 The rims cost $700. t + 700 = 2300 t = 1600 The tires cost $1600 which is $200 more than twice the cost of the $700 rims.